# Freestyle Sketch-in-the-Loop Image Segmentation

### v1

@@ Abstract
In this paper, we expand the domain of sketch research into the field of image segmentation, aiming to establish freehand sketches as a query modality for subjective image segmentation. Our innovative approach introduces a “sketch-in-the-loop” image segmentation framework, enabling the segmentation of visual concepts partially, completely, or in groupings – a truly “freestyle” approach – without the need for a purpose-made dataset (i.e., mask-free). This framework capitalises on the synergy between sketch-based image retrieval (SBIR) models and large-scale pre-trained models (CLIP or DINOv2). The former provides an effective training signal, while fine-tuned versions of the latter execute the subjective segmentation. Additionally, our purpose-made augmentation strategy enhances the versatility of our sketch-guided mask generation, allowing segmentation at multiple granularity levels. Extensive evaluations across diverse benchmark datasets underscore the superior performance of our method in comparison to existing approaches across various evaluation scenarios.
@@ 1 Introduction
This paper delves into the realm of image segmentation, but with a unique twist – the incorporation of freehand sketches. Our research vision is to explore the impact of fine-grained subjectivity inherently captured by a sketch [30, 19], on image segmentation. At its core, our aim is to devise a subjective image segmentation framework that empowers users to freely segment any object, be it a partial segment, a complete entity, or a group of interconnected elements – essentially, enabling image segmentation in a “freestyle” manner.
Image segmentation has witnessed significant evolution [14, 39] over time, and we are committed to keeping pace with the current capabilities of image segmentation systems – we also want to operate in an open-vocabulary [57] manner and harness the potential of the latest foundation models. We too fully acknowledge prior segmentation methods that inject subjectivity through textual inputs [85]. The central argument of this paper however revolves around the premise that sketches offer a more precise form of subjectivity – a point convincingly illustrated in LABEL:fig:teaser (bottom).
To elucidate why sketches serve as a superior modality for capturing subjectivity, we direct readers to the extensive sketch literature [10, 19, 86, 25, 74]. In summary, the insights drawn from this large body of work are two-fold: (i) sketches excel in their ability to convey intricate visual details that may be cumbersome or ambiguous when expressed through text [74]. (ii) sketches exhibit a smaller semantic gap to photos, as both share the same visual modality, albeit in an abstract form [30]. This inherent relationship is notably featured in the extensively studied domain of SBIR [20] and its fine-grained variations [7] – a challenge that we also employ as a guiding critic in our training.
The specifics of our vision present significant challenges. First and foremost, we need to operate without any sketch-mask pairing data. The collection of such a dataset is simply unfeasible, especially considering the current status quo within the sketch community still concerns gathering enough sketch-photo pairs [58, 73], let alone pixel-perfect masks. This challenge is further exacerbated by our open-vocabulary setting, which essentially renders prior arts inapplicable [31]. Secondly, akin to most prior work in the sketch literature [86, 72, 6, 41], we face a substantial gap between the sketch and photo domains.
We commence by fixating on the ultimate goal: given a sketch, compute corresponding segmentation masks for all applicable gallery photos. Surprisingly, it turns out that all we need to compute this correspondence is the strategic utilisation of (i) a frozen Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) model and (ii) a pre-trained foundation model (e.g., CLIP [68], DINOv2 [60]). The former provides a solid foundation for fine-grained sketch-photo correspondence and also serves as a training signal to fine-tune the latter. Meanwhile, the fine-tuned version of the latter predicts the actual mask for the photo using its corresponding sketch, thereby eliminating the need for manually annotated masks (i.e., mask-free). It follows that these two components are intricately linked by this very hypothesis: a sketch and its matching background-masked image pair (and foreground highlighted) will sit closely within the fixed FG-SBIR joint embedding [72]. This, we hope, is intuitive – the more highlighted (using the computed mask) the foreground object is, the easier it is for its matching sketch to locate. It is through this iterative training process that the segmentation model continually hones its performance.
In generating sketch-guided masks, we use pre-trained DINOv2 [60] for the fine-grained setup (CLIP [68] for category-level). It extracts a spatial feature map from a photo and a global feature vector from a sketch. A cosine similarity computation, followed by bilinear upscaling and differentiable thresholding, yields a pseudo mask used for generating the background-masked image corresponding to the given sketch. To prevent all-white mask predictions and overfitting, a mask-regularisation loss is introduced. Additionally, for part-level segmentation, we employ sketch-partitioning, randomly fragmenting binary sketches into orthogonal splits, enabling model-prediction from a specific split, for more precise part-level segmentation.
In summary, our contributions are (i) we address the challenge of subjective segmentation by introducing a novel approach that enables mask-free image segmentation through the interpretive power of sketches. (ii) we support multi-granularity segmentation, providing users with the flexibility to segment visual concepts at various levels, encompassing entire entities, specific parts, or interconnected groups. (iii) we establish an effective synergy between sketch-based image retrieval models and large-scale pre-trained models. (iv) we propose a purpose-made augmentation strategy that enhances the framework’s versatility, enabling segmentation at multiple granularity levels.
@@ 2 Related Works
Sketch for Vision Tasks.Freehand sketches has long established its significance for several visual understanding tasks [8, 86, 72, 19, 41, 10]. Following its success in $2D$ and $3D$ image retrieval [86, 72, 24, 54, 53], sketch as a vision-modality has been used in image generation [41], saliency detection [10], inpainting [87], shape-modelling [56], augmented reality [54], medical image analysis [40], image editing [90, 47], object detection [19], representation learning [22], class-incremental learning [8], etc. To our best knowledge, Sketch-a-Segmenter [31] is the only existing work dealing with sketch-guided segmentation. This method [31] however requires fully-annotated pixel-level masks for training [31], unlike ours. We thus here aim to forward sketch research in another interesting direction by utilising its potential as a weak label for segmentation.
@@ 4 Proposed Methodology
Overview.  Given an image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, state-of-the-art supervised image segmentation models [49, 14, 15] aim to generate a segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times(c+1)}$ via pixel-level classification across a pre-defined set of $c$ classes and a background ($+1$). On the contrary, given a query sketch $\mathcal{S}$, our method generates a binary segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$, depicting the pixel-locations, wherever the queried-concept appears in any candidate image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, with pixels belonging to that concept marked as $1$ and all the rest as 00. Learning segmentation frameworks via typical supervised training setup however, is impractical given the laborious process of collecting human-annotated ground truth segmentation masks. We therefore aim to utilise the existing sketch-photo dataset [73] to devise a weakly-supervised mask-free image segmenter. Although existing weak supervision signals (query) like tags [2], caption [57], cue points [5], etc. might suffice for category-level vision tasks, they lack in fine-grained cues unlike sketches [10, 19]. Inspired by sketch’s success in multiple fine-grained vision tasks (e.g., retrieval [72], detection [19], generation [41], etc.), we aim to establish sketch as a worthy query-modality for mask-free weakly-supervised image segmentation, both at category-level and fine-grained setup. Furthermore, we also extend our method to part-level segmentation, using the same dataset.
@@ 4.2 Baseline Weakly Supervised Segmentation
In our setup, we consider the available sketches ($\mathcal{S}$) and photos ($\mathcal{P}$) from $\mathcal{D}_{\mathrm{cat}}$= $\{\mathcal{S}_{i},\mathcal{P}_{i}\}_{i=1}^{M}$ and $\mathcal{D}_{\mathrm{fine}}$=$\{\mathcal{S}_{j},\mathcal{P}_{j}\}_{j=1}^{N}$ with category-level [48] and fine-grained [73] association between them respectively (cf. Sec. 3). We design our baseline sketch-based image segmenter taking inspiration from text-based weakly-supervised segmentation framework [57] that employs patch-wise InfoNCE loss between CLIP’s global textual feature and patch-wise visual features [57] for better text-photo region-wise alignment (instead of using global feature for both as in CLIP training [68]) required for dense pixel-level segmentation task [57]. We however exclude the textual encoder and employ CLIP’s visual encoder to extract patch-wise photo features and global query sketch feature to apply the same loss.
@@ 4.3 Sketch-guided Mask Generation
$\small\mathcal{C}(u,v)=\frac{\langle\mathbf{s},\mathbf{p}_{s}(u,v)\rangle}{||\mathbf{s}||~{}||\mathbf{p}_{s}(u,v)||}~{}~{}~{}~{}u=[1,..,h^{\prime}],v=[1,..,w^{\prime}]\vspace{-0.1cm}$ (2)
@@ 4.5 Further Extension to Part-level Segmentation
Part-level segmentation aims to restrict the model to segment only the region corresponding to the part-level sketch query (e.g., “head” of a camel). Unfortunately, common sketch-photo datasets [73] neither hold part-level sketches, nor part-level masks. We therefore aim to model this part-level behaviour via a simple sketch augmentation trick where we synthetically split the input sketch into multiple non-overlapping portions. Specifically, we first calculate the centroid of the query sketch from its vector coordinates. Now, at every pass, we draw a straight line from the centroid with a random slope $m_{1}$. Based on the number of non-overlapping portions $n$, we calculate the slope values ($m_{2},m_{3},\cdots,m_{n}$) for the next lines as $m_{n}$=$m_{n-1}+\frac{360}{n}$. These $n$ straight lines segment the sketch into $n$ orthogonal non-overlapping parts (Fig. 4). With $n$=$2$ (decided empirically), we segment the query sketch $\mathcal{S}$ into two non-overlapping parts $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$. Using sketch-guided mask generation procedure (Sec. 4.3), we generate the corresponding masks $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$. Ideally, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ should only segment their corresponding regions from the candidate images in a non-overlapping manner, which therefore incurs a penalisation if $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ overlap. To do so we combine $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ via bit-wise XOR to get the final mask as $\mathcal{M}$=$\mathcal{M}_{A}\oplus\mathcal{M}_{B}$, where the common foreground regions should be nullified. Consequently, the same regions will be suppressed in the masked candidate image $\mathcal{P}_{\mathrm{mask}}$=$\mathcal{P}\times{\mathcal{M}}$ as well. Accordingly, $\mathcal{L}_{\mathrm{SBIR}}$ would impose a high loss, leading to better part-specific segmentation.
@@ 5 Experiments
Implementation Details.  We use a pre-trained ViT-L/14 CLIP [68] and ViT-L/14 DINOv2 [60] backbone for category-level and fine-grained segmentation models respectively. We fine-tune the $\mathtt{LayerNorm}$ layers of the CLIP visual encoder with a learning rate of $10^{-5}$, keeping the rest of the parameters frozen. The DINOv2 encoder is fine-tuned with a low learning rate of $10^{-6}$. We train the model for $200$ epochs, in an NVIDIA RTX 2080-Ti GPU using AdamW [50] optimiser and a batch size of $16$. All $\lambda$ values are set to $1$, empirically. We use pre-trained baseline SBIR and FG-SBIR models from [71].
@@ 5.2 Ablation on Design
[ii] Does Fine-tuning entire CLIP help?  Existing literature on sketch-based vision tasks [74, 72] point towards fine-tuning the entire CLIP vision encoder to adapt it to their respective tasks [74, 72]. Doing so for training our category-level segmentation model however, significantly drops the final seen/unseen class mIoU (CLIP Fine-tuning result in Tab. 2). We hypothesise that fine-tuning all parameters of the vision encoder distorts the pre-trained knowledge of CLIP, thus sacrificing its generalisation potential.
[iii] Effect of $\mathcal{L}_{\mathrm{InfoNCE}}$.  We posit that the InfoNCE loss aids in proper alignment of sketch-photo features to generate accurate patch-wise correlation maps. Hence, a drop of $12.69~{}(17.48)\%$ seen-class mIoU in w/o $\mathcal{L}_{\mathrm{InfoNCE}}$ (Tab. 2), for category-level (fine-grained) setup, verifies its importance in learning patch-wise sketch-photo region alignment.
@@ 6 Conclusion
Here, we explore the potential of freehand abstract sketches for subjective image segmentation. Leveraging large-scale pre-trained visual encoders and SBIR models, we devise a mask-free image segmenter that allows for segmentation at multiple granularity levels (i.e., category-level and fine-grained). Furthermore, using the smart augmentation trick of sketch-partitioning, we extend our method to part-level segmentation. Extensive qualitative and quantitative results depict the efficacy of our method. We hope that this study paves the path for more sketch-based vision tasks in future.

### v2

# SketchYourSeg: Mask-Free Subjective Image Segmentation via Freehand Sketches

@@ Abstract
We introduce SketchYourSeg, a novel framework that establishes freehand sketches as a powerful query modality for subjective image segmentation across entire galleries through a single exemplar sketch. Unlike text prompts that struggle with spatial specificity or interactive methods confined to single-image operations, sketches naturally combine semantic intent with structural precision. This unique dual encoding enables precise visual disambiguation for segmentation tasks where text descriptions would be cumbersome or ambiguous – such as distinguishing between visually similar instances, specifying exact part boundaries, or indicating spatial relationships in composed concepts. Our approach addresses three fundamental challenges: (i) eliminating the need for pixel-perfect annotation masks during training with a mask-free framework; (ii) creating a synergistic relationship between sketch-based image retrieval (SBIR) models and foundation models (CLIP/DINOv2) where the former provides training signals while the latter generates masks; and (iii) enabling multi-granular segmentation capabilities through purpose-made sketch augmentation strategies. Our extensive evaluations demonstrate superior performance over existing approaches across diverse benchmarks, establishing a new paradigm for user-guided image segmentation that balances precision with efficiency.
@@ 1 Introduction
This paper introduces a novel approach to image segmentation that leverages freehand sketches for subjective segmentation across entire galleries with a single exemplar sketch. Our method, SketchYourSeg, enables users to express personalised segmentation intent through simple sketches rather than complex text descriptions or repetitive interactive inputs, whether targeting whole objects, specific parts, or multiple related elements.
Image segmentation has evolved substantially [36], with current systems increasingly leveraging foundation models in open-vocabulary settings [55]. However, a critical question remains: how can users effectively communicate their subjective segmentation intent while efficiently operating across multiple images? While text prompts [85] and interactive methods [16, 26] have gained popularity, they present fundamental limitations for precisely specifying segmentation targets, especially across entire image sets.
To address this challenge, we investigate freehand sketches as a query modality for subjective segmentation. A key question arises: why use sketches when alternatives like text prompts or interactive scribbles exist? Research evidence demonstrates that sketches occupy a unique middle ground in the specificity-effort tradeoff curve. Text prompts, while easy to produce and capable of gallery-wide operations, struggle with spatial and structural specificity for fine-grained tasks [39, 56, 17, 75, 9]. For instance, precisely describing “the zebra with distinctive stripes that curve inward at the shoulder” through text alone presents significant challenges that sketches can naturally overcome. Interactive methods using clicks or scribbles, while offering precise spatial control, operate only on one image at a time and typically require multiple interactions per image [80]. This fundamental limitation makes them impractical for gallery-wide operations where users need to segment the same concept across multiple images simultaneously.
Sketches, by contrast, naturally combine semantic intent with structural specificity in a single interaction [40]. This dual encoding enables our approach to segment visual concepts across entire image galleries through a single exemplar sketch – a capability neither text nor interactive methods can match. The sketch serves as both a semantic concept descriptor (similar to text) and a structural template (similar to scribbles), making it uniquely suited for fine-grained segmentation tasks where spatial relationships and distinctive visual features play a crucial role.
Our approach confronts two significant technical challenges. First, we must operate without sketch-mask pairing data, as collecting such a dataset is infeasible given the sketch community’s ongoing struggle to gather even basic sketch-photo pairs [56, 71]. Second, we must bridge the substantial domain gap between abstract sketches and photos [86, 70] at the pixel level.
We address these challenges through a novel mask-free approach that strategically utilizes: (1) a frozen Fine-Grained Sketch-Based Image Retrieval (FG-SBIR) model, and (2) a pre-trained foundation model (CLIP [66] or DINOv2 [58]). These components are linked by our central hypothesis: a sketch and its matching background-masked image will occupy proximate positions within a fixed FG-SBIR embedding space [70]. This creates a self-supervised loop – the better our segmentation mask highlights the relevant content, the closer the masked image and sketch will be in the embedding space, providing a natural training signal without manual annotations.
For mask generation, we extract spatial feature maps from photos and global feature vectors from sketches. Using cosine similarity, bilinear upscaling, and differentiable thresholding, we generate pseudo masks that highlight regions corresponding to the sketch query. To prevent trivial solutions and enable multi-granular segmentation, we introduce mask regularisation techniques and a novel sketch-partitioning strategy that enables part-level segmentation without additional training data.
Our primary contributions are: (i) A mask-free training paradigm that eliminates the need for pixel-perfect annotations while still achieving high-quality segmentation results. (ii) A synergistic framework that leverages sketch-based image retrieval models as critics to guide foundation models in generating accurate segmentation masks. (iii) Multi-granular segmentation capabilities that leverage sketches’ unique structural encoding to precisely specify visual intent that text cannot effectively communicate – from shape-specific distinctions between similar instances to exact part boundaries and spatial relationships between objects – all within a single framework. Empirical validation through extensive benchmark evaluations, demonstrating superior performance over existing approaches for both seen and unseen categories.
@@ 2 Related Works
Sketch for Vision Tasks.Freehand sketches has long established its significance for several visual understanding tasks [7, 86, 70, 18, 38, 9]. Following its success in $2D$ and $3D$ image retrieval [86, 70, 23, 52, 51], sketch as a vision-modality has been used in image generation [38], saliency detection [9], inpainting [87], shape-modelling [54], augmented reality [52], medical image analysis [37], image editing [90, 46], object detection [18], representation learning [21], class-incremental learning [7], etc. To our best knowledge, Sketch-a-Segmenter [28] is the only existing work dealing with sketch-guided segmentation.
@@ 4 Proposed Methodology
Overview.  Given an image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, state-of-the-art supervised image segmentation models [48, 13, 14] aim to generate a segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times(c+1)}$ via pixel-level classification across a pre-defined set of $c$ classes and a background ($+1$). On the contrary, given a query sketch $\mathcal{S}$, our method generates a binary segmentation mask $\mathcal{M}\in\mathbb{R}^{h\times w\times 1}$, depicting the pixel-locations, wherever the queried-concept appears in any candidate image $\mathcal{I}\in\mathbb{R}^{h\times w\times 3}$, with pixels belonging to that concept marked as $1$ and all the rest as 00. Learning segmentation frameworks via typical supervised training setup however, is impractical given the laborious process of collecting human-annotated ground truth segmentation masks. We therefore aim to utilise the existing sketch-photo dataset [71] to devise a weakly-supervised mask-free image segmenter. Although existing weak supervision signals (query) like tags [2], caption [55], cue points [5], etc. might suffice for category-level vision tasks, they lack in fine-grained cues unlike sketches [9, 18]. Inspired by sketch’s success in multiple fine-grained vision tasks (e.g., retrieval [70], detection [18], generation [38], etc.), we aim to establish sketch as a worthy query-modality for mask-free weakly-supervised image segmentation, at category-level, fine-grained, and part-level segmentation setup.
@@ 4.2 Baseline Weakly Supervised Segmentation
We consider the available sketches ($\mathcal{S}$) and photos ($\mathcal{P}$) from $\mathcal{D}_{\mathrm{cat}}$=$\{\mathcal{S}_{i},\mathcal{P}_{i}\}_{i=1}^{M}$ and $\mathcal{D}_{\mathrm{fine}}$=$\{\mathcal{S}_{j},\mathcal{P}_{j}\}_{j=1}^{N}$ with category-level and fine-grained association between them respectively. We design the baseline sketch-based image segmenter taking inspiration from text-based weakly-supervised segmentation method [55] that uses patch-wise InfoNCE loss between CLIP’s global textual feature and patch-wise visual features [55] for better text-photo region-wise alignment (instead of using global feature for both as in CLIP training) required for dense pixel-level segmentation task. We however exclude the textual encoder and employ CLIP’s visual encoder to extract patch-wise photo features and global query sketch feature to apply the same loss.
@@ 4.3 Sketch-guided Mask Generation
$\mathcal{C}(u,v)=\frac{\langle\mathbf{s},\mathbf{p}_{s}(u,v)\rangle}{||\mathbf{s}||~{}||\mathbf{p}_{s}(u,v)||}~{}~{}~{}~{}u=[1,..,h^{\prime}],v=[1,..,w^{\prime}]\vspace{-0.1cm}$ (2)
@@ 4.5 Further Extension to Part-level Segmentation
Part-level segmentation aims to restrict the model to segment only the region corresponding to the part-level sketch query (e.g., “head” of a camel). Unfortunately, common sketch-photo datasets [71] neither hold part-level sketches, nor part-level masks. We therefore aim to model this part-level behaviour via a simple sketch augmentation trick where we synthetically split the input sketch into multiple non-overlapping portions. Specifically, we first calculate the centroid of the query sketch from its vector coordinates. Now, at every pass, we draw a straight line from the centroid with a random slope $m_{1}$. Based on the number of non-overlapping portions $n$, we calculate the slope values ($m_{2},m_{3},\cdots,m_{n}$) for the next lines as $m_{n}$=$m_{n-1}+\frac{360}{n}$. These $n$ straight lines segment the sketch into $n$ orthogonal non-overlapping parts (Fig. 4).
@@
With $n$=$2$ (decided empirically), we segment the query sketch $\mathcal{S}$ into two non-overlapping parts $\mathcal{S}_{A}$ and $\mathcal{S}_{B}$. Using sketch-guided mask generation procedure (Sec. 4.3), we generate the corresponding masks $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$. Ideally, $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ should only segment their corresponding regions from the candidate images in a non-overlapping manner, which therefore incurs a penalisation if $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ overlap. To do so we combine $\mathcal{M}_{A}$ and $\mathcal{M}_{B}$ via bit-wise XOR to get the final mask as $\mathcal{M}$=$\mathcal{M}_{A}\oplus\mathcal{M}_{B}$, where the common foreground regions should be nullified. Consequently, the same regions will be suppressed in the masked candidate image $\mathcal{P}_{\mathrm{mask}}$=$\mathcal{P}\times{\mathcal{M}}$ as well. Accordingly, $\mathcal{L}_{\mathrm{SBIR}}$ would impose a high loss, leading to better part-specific segmentation.
@@ 5 Experiments
Implementation Details.  We use a pre-trained ViT-L/14 CLIP [66] and ViT-L/14 DINOv2 [58] backbone for category-level and fine-grained segmentation models respectively. We fine-tune the $\mathtt{LayerNorm}$ layers of the CLIP visual encoder with a learning rate of $10^{-5}$, keeping the rest of the parameters frozen. The DINOv2 encoder is fine-tuned with a low learning rate of $10^{-6}$. We train the model for $200$ epochs using AdamW [49] optimiser and a batch size of $16$. All $\lambda$ values are set to $1$, empirically. We use pre-trained baseline SBIR and FG-SBIR models from [69].
@@ 5.2 Ablation on Design
[ii] Does Fine-tuning entire CLIP help?  Existing literature on sketch-based vision tasks [72, 70] point towards fine-tuning the entire CLIP vision encoder to adapt it to their respective tasks [72, 70]. Doing so for training our category-level segmentation model however, significantly drops the final seen/unseen class mIoU (CLIP Fine-tuning result in Tab. 2). We posit that fine-tuning all parameters of the vision encoder distorts the pre-trained knowledge of CLIP, thus sacrificing its generalisation potential.
[iii] Effect of $\mathcal{L}_{\mathrm{InfoNCE}}$.  We posit that the InfoNCE loss aids in proper alignment of sketch-photo features to generate accurate patch-wise correlation maps. Hence, a drop of $12.69~{}(17.48)\%$ seen-class mIoU in w/o $\mathcal{L}_{\mathrm{InfoNCE}}$, for category-level (fine-grained) setup, verifies its importance in learning patch-wise sketch-photo region alignment.
@@ 6 Conclusion
Here, we explore the potential of freehand abstract sketches for subjective image segmentation. Leveraging large-scale pre-trained visual encoders and SBIR models, we devise a mask-free image segmenter that allows for segmentation at multiple granularity levels (i.e., category-level and fine-grained). Furthermore, using the smart augmentation trick of sketch-partitioning, we extend our method to part-level segmentation. Extensive qualitative and quantitative results depict the efficacy of our method.